OpenStudy (anonymous):

find the length of molding needed to put around a circular table that is 3.8 ft in diameter. use 3.14 for pi.

OpenStudy (ghazi):

required length = perimeter of circle = 2*3.14* (3.8/2) radius = 3.8/2= 1.9